Show the bulk modulus of water is 2.1 x 10 N/m if the speed of sound is 1450 m/s. (p=1.0 x 10
1
Expert's answer
2021-10-10T18:11:49-0400

"v=\sqrt{\frac E{\rho}},\implies"

"E=v^2\rho,"

"E=1450^2\cdot 1.0\cdot 10^3=2.1\cdot 10^9~\frac N{m^2}."